40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Entry  Tue Dec 10 17:03:41 2019, anchal, DailyProgress, Laser, Laser Settings back to defaults 
    Reply  Wed Dec 18 14:04:43 2019, anchal, DailyProgress, Laser, South Laser Power Dropped. SouthLaserPowerDropped.pdf
Message ID: 2501     Entry time: Wed Dec 18 14:04:43 2019     In reply to: 2492
Author: anchal 
Type: DailyProgress 
Category: Laser 
Subject: South Laser Power Dropped. 

After the last calibration, South Laser power at Laser head determined by its internal photodiode and show as setting PWR today (noticed today) dropped to 80 mW with none of the other settings changed. The diode current is still the same, so this could be one of the following two things:
1) The internal circuit to measure laser power at head went wrong.

2) Something is wrong with the laser crystal.

The laser power as seen by the reflection photodiode at South PMC captured a glitch on Dec 17th, 2019 8:30 am (might be 7:30 am in PST). The glitch shows that PMC went out of lock due to it but then returned to a laser power lower than before. The voltage level decreased from 2.72 V to 2.3 V after lock, which is equivalent to a drop of ~15.4%. The laser head power level according to PWR monitor dropped from 101 mW to 80 mW, which is equivalent to ~20.8%. While these don't match, PMC reflection is also not a true measure of power level. But I am sure this power has actually reduced and even if the laser head power meter is off a little bit, we have witnessed a big drop in power.

Restarting the laser (Put it on Standby and switch it back on) didn't change the power level.

Restarting the laser with turning the key off and on also didn't change the power level.

Even though 80 mW is enough for our experiment, this sudden decline in power shows there is something happening with the laser that we do not understand.


Attachment 1: SouthLaserPowerDropped.pdf  27 kB  | Hide | Hide all
ELOG V3.1.3-